A sample of size 60 is taken from a population of chocolate balls. The weights of chocolate balls are normally distributed with a mean of 40 g and a standard deviation of 8 g. a. Find the parameters for this distribution of the sample mean. b. Find the probability that the sample mean is between 38 g and 42 


1
Expert's answer
2021-06-24T07:27:34-0400

Let "\bar{X}=" the sample mean: "\bar{X}\sim N(\mu, \sigma^2/n)."

a.


"\mu_{\bar{X}}=\mu=40\ g, \sigma_{\bar{X}}=\sigma/\sqrt{n}=8/\sqrt{60}\ g"

b.


"P(38<\bar{X}<42)=P(\bar{X}<42)-P(\bar{X}\leq38)"

"=P(Z<\dfrac{42-40}{8/\sqrt{60}})-P(Z\leq\dfrac{38-40}{8/\sqrt{60}})"

"\approx P(Z<1.93649)-P(Z\leq-1.93649)"


"\approx 0.973596-0.026404\approx0.9472"

The probability that the sample mean is between 38 g and 42 g is 0.9472.